OpenStudy (anonymous):

Determine the positive integer values of k for which the following polynomial is factorable over the integers. (Enter your answers as a comma-separated list.) z2 + 6z + k

OpenStudy (anonymous):

9 would work because then it would factor as \[x^2+6x+9=(x+3)^2\] 5 would work as well because you would get \[x^2+6x+5=(x+1)(x+5)\] and i believe